# Classicized dynamics and initial conditions in field theories with fakeons

### v1

@@ Abstract
Theories with purely virtual particles (fakeons) do not possess a classical action in the strict sense, but rather a classicized one, obtained by integrating out the fake particles at tree level. Although this procedure generates nonlocal interactions, we show that the resulting classicized equations of motion are not burdened with the need to specify infinitely many initial conditions. The reason is the inherent link between the fakeonic system and the parent higher-derivative local system: the solution space of the former is an appropriate subspace of solutions of the latter. A somewhat unexpected proviso is that, in order to avoid overcounting, the fakeon prescription must be obtained as a limit or special case of a more generic prescription. Ultimately, the number of degrees of freedom matches physical expectations, the extra ones (ghosts or otherwise) being removed by rendering them purely virtual. We illustrate the counting in simple linear solvable models and provide the general proof. Along similar lines, we analyze Dirac’s removal of runaway solutions in classical electrodynamics.
@@ 1 Introduction
Fakeons offer a relatively simple resolution of the tension between the requirement of renormalizability and the violation of unitarity induced by higher-derivative operators. This goal can be achieved in four different but equivalent ways. The first two methods rely on the Anselmi–Piva (AP) or fakeon prescription, which can be implemented either 1) by combining the Lee–Wick prescription Lee:1969fy ; Lee:1970iw ; Lee:1969zze ; Cutkosky:1969fq ; Nakanishi:1971jj for loop-energy integrals with a suitable deformation of the spatial-momentum integration domain Anselmi:2017yux ; Anselmi:2017lia ; Anselmi:2018kgz , or 2) by averaging the analytically continued Euclidean amplitude around branch points involving the would-be fakeons Anselmi:2017yux ; Anselmi:2017lia ; Anselmi:2018kgz . Alternatively, one can 3) introduce a new type of diagrammar Anselmi:2021hab , or 4) use specially defined non-time-ordered correlation functions Anselmi:2022qor .
@@
Several aspects of the classicization deserve emphasis. First, the classicized equations of motion are approximate, because the correction terms are computed perturbatively in the couplings. Finding the fully resummed classicized Lagrangian can be highly nontrivial, but it is expected: purely virtual systems are intrinsically quantum and classicization carries typical quantum features (such as perturbativity) down to the classical limit.
@@
We show that fakeonic nonlocal systems generated by classicization are immune to this difficulty, due to their link to the parent local higher-derivative system (the one associated with the interim action), which involves only finitely many time derivatives. The equations of the interim Lagrangian are recovered by acting on the classicized equations with a suitable differential operator. Hence, the solutions of the classicized equations form a subspace of those that solve the interim system, and involve only a finite number of arbitrary constants. We further show that the degrees of freedom of the fakeonic system match those of its physical subsector.
@@ 3.2 Classicization of higher-derivative Lagrangians with a source
For example, the naturally accelerated motion ($F=$ constant) is unmodified: $\ddot{q}(t)=F$.
@@ 4 Initial conditions and number of degrees of freedom
$\Box\varphi+m^{2}\varphi=\frac{g^{2}}{2}\varphi\frac{M^{2}}{(\Box+m^{2})^{2}+M^{4}}\varphi^{2},$
@@
$\Box\varphi_{0}+m^{2}\varphi_{0}=0,\qquad\Box\varphi_{n}+m^{2}\varphi_{n}=\frac{1}{2}\sum_{i+j+k=n-1}\varphi_{i}\frac{M^{2}}{(\Box+m^{2})^{2}+M^{4}}\varphi_{j}\varphi_{k}\,,\qquad n>0.$
@@ 4.1.1 Dirac’s removal of runaway solutions
In this way, the higher-derivative local system (4.40), which has one runaway solution, is converted into an inequivalent nonlocal system, (4.41), which lacks the runaway solution (as we show below in detail) and is fixed by one initial condition less. Moreover, the latter violates microcausality, since the right-hand side requires knowledge of the external force in a “little bit of future” (specifically, an amount of future of order $\tau$) to predict the future beyond that.
@@ 4.1.3 Tachyonic fakeon two-derivative model with repulsive force
What just said teaches us that if we want to “fakeonize” an operator, we have to work in a generic environment (i.e., use the most general Green function that inverts the operator), where the right number of initial conditions follows straightforwardly. Particular cases must be reached as limits.
@@ 4.2 General proof
$G(x,b_{j})=G_{\textrm{AP}}(x)+\int\frac{{\rm d}^{3}\bm{p}}{(2\pi)^{3}}\,{\rm e}^{i\bm{p}\cdot\bm{x}}\left[b_{+}(\bm{p})\,{\rm e}^{-ix^{0}\Omega}+b_{-}(\bm{p})\,{\rm e}^{ix^{0}\Omega}\right]+\textrm{c.c.},$
@@
A quick way to see this is as follows. Insert the functions $\varphi(x,a_{i})$ into the right-hand side of (4.59). We have no guarantee that the result is zero, but we know that that is a zero mode of the operator $(\Box^{2}+M^{4})/M^{2}$, because of (4.58). Let us denote it by $\psi_{0}$. The kernel of $(\Box^{2}+M^{4})/M^{2}$ is a four-dimensional space. Generically, we need to impose four relations among the parameters $a_{i}$ and $b_{j}$ to make $\psi_{0}$ disappear.

### v2

@@ Abstract
Theories with purely virtual particles (fakeons) do not possess a classical action in the strict sense, but rather a classicized one, obtained by integrating out the fake particles at tree level. Although this procedure generates nonlocal interactions, we show that the resulting classicized equations of motion are not burdened with the need to specify infinitely many initial conditions. The reason is the inherent link between the fakeonic system and the parent higher-derivative local system: the solution space of the former is an appropriate subspace of solutions of the latter. A somewhat unexpected proviso is that, in order to avoid overcounting, the fakeon prescription must be obtained as a limit or special case of a more generic prescription. Ultimately, the set of degrees of freedom matches physical expectations, the extra ones (ghosts or otherwise) being removed by rendering them purely virtual. We illustrate the counting in simple linear solvable models and provide the general proof. Along similar lines, we analyze Dirac’s removal of runaway solutions in classical electrodynamics.
@@ 1 Introduction
Fakeons offer a relatively simple resolution of the tension between the requirement of renormalizability and the violation of unitarity induced by higher-derivative operators. This goal can be achieved in four different but equivalent ways. The first two methods rely on the Anselmi–Piva (AP) or fakeon prescription, which can be implemented either 1) by combining the Lee–Wick prescription Lee:1969fy ; Lee:1970iw ; Lee:1969zze ; Cutkosky:1969fq ; Nakanishi:1971jj for loop-energy integrals with a suitable deformation of the spatial-momentum integration domain Anselmi:2017yux ; Anselmi:2017lia ; Anselmi:2018kgz , or 2) by averaging the analytically continued Euclidean amplitude around branch points involving the would-be fakeons Anselmi:2017yux ; Anselmi:2017lia ; Anselmi:2018kgz . Alternatively, one may 3) introduce a new type of diagrammar Anselmi:2021hab , or 4) employ specially defined non-time-ordered correlation functions Anselmi:2022qor .
The standard Cutkosky identities, which express the unitarity equation $S^{\dagger}S=1$ in diagrammatic form, can be generalized to theories involving fakeons, for arbitrary diagrams Anselmi:2021hab ; Anselmi:2022qor . In brief Anselmi:2021hab , one first refines the identities in theories containing only physical particles, recognizing that they decompose into a set of independent “spectral optical identities,” each holding at fixed spatial momenta and involving the same threshold. It then follows that fakeonizing a particle simply corresponds to replacing the standard Feynman diagrams with combinations of diagrams that $i$) eliminate the spectral optical identities associated with thresholds involving at least one fakeon, and ii)ii) leave all remaining identities unchanged.
@@
Several aspects of the classicization deserve emphasis. First, the classicized equations of motion are approximate, because the correction terms are computed perturbatively in the couplings. Finding the fully resummed classicized Lagrangian can be highly nontrivial, but this is expected: purely virtual systems are intrinsically quantum and classicization carries typical quantum features (such as perturbativity) down to the classical limit.
@@
We show that fakeonic nonlocal systems generated by classicization are immune to this difficulty, due to their link to the parent local higher-derivative system (the one associated with the interim action), which involves only finitely many time derivatives. The equations of the interim Lagrangian are recovered by acting on the classicized equations with a suitable differential operator. Hence, the solutions of the classicized equations form a subspace of those that solve the interim system, and involve only a finite number of arbitrary constants. We further show that the degrees of freedom of the fakeonic system match those of its expected physical subsector.
@@ 3 Classicization
Note that the classicized action is different from the more familiar one-loop effective action. The classicized action is obtained by integrating out only the fakeons (treating the other fields as external) and just keeping the tree diagrams, while the one-loop effective action is obtained by integrating on all the fields up to the one-loop diagrams. A possible source of concern is the infrared problems arising in the classical limit when massless particles are integrated out. The reason why fakeons are not affected by this issue is that they must be massive, because the reciprocal $1/m_{\chi}$ of their mass $m_{\chi}$ measures the violation of (micro)causality. Massless (or too light) fakeons are unacceptable, since they trigger a violation of causality, contrary to evidence at our scales. In quadratic gravity, for example (see subsection 3.3), the classicization amounts to just integrating out the heavy spin-two gravifakeon $\chi_{\mu\nu}$ at the tree level, treating the graviton as an external field.
@@ 3.2 Classicization of higher-derivative Lagrangians with a source
For example, the uniformly accelerated motion ($F=$ constant) is unmodified: $\ddot{q}(t)=F$.
@@ 4 Initial conditions and number of degrees of freedom
$\Box\varphi+m^{2}\varphi=\frac{g^{2}}{2}\varphi\left.\frac{M^{2}}{(\Box+m^{2})^{2}+M^{4}}\right|_{\textrm{f}}\varphi^{2},$
@@
$\Box\varphi_{0}+m^{2}\varphi_{0}=0,\qquad\Box\varphi_{n}+m^{2}\varphi_{n}=\frac{1}{2}\sum_{i+j+k=n-1}\varphi_{i}\left.\frac{M^{2}}{(\Box+m^{2})^{2}+M^{4}}\right|_{\textrm{f}}\varphi_{j}\varphi_{k}\,,\qquad n>0.$
@@ 4.1.1 Dirac’s removal of runaway solutions
In this way, the higher-derivative local system (4.40), which has one runaway solution, is converted into an inequivalent nonlocal system, (4.41), which lacks the runaway solution (as we show below in detail) and is fixed by one initial condition less. Moreover, the latter violates microcausality, since the right-hand side requires knowledge of the external force in a “little bit of future” (specifically, an amount of future of order $\tau$) beyond the future we want to predict.
@@ 4.1.3 Tachyonic fakeon two-derivative model with repulsive force
What just said teaches us that if we want to “fakeonize” an operator, we have to work in a generic environment (i.e., use the most general Green function that inverts the operator), where the right number of initial conditions follows straightforwardly. Particular cases, such as the fakeon one, must be reached as limits.
@@ 4.2 General proof
$G(x,b_{j})=G_{\textrm{AP}}(x)+\left\{\int\frac{{\rm d}^{3}\bm{p}}{(2\pi)^{3}}\,{\rm e}^{i\bm{p}\cdot\bm{x}}\left[b_{+}(\bm{p})\,{\rm e}^{-ix^{0}\Omega}+b_{-}(\bm{p})\,{\rm e}^{ix^{0}\Omega}\right]+\textrm{c.c.}\right\},$
@@
A quick way to see this is as follows. Insert the functions $\varphi(x,a_{i})$ into the right-hand side of (4.59). We have no guarantee that the result is zero, but we know that what we obtain is a zero mode of the operator $(\Box^{2}+M^{4})/M^{2}$, because of (4.58). Let us denote it by $\psi_{0}$. The kernel of $(\Box^{2}+M^{4})/M^{2}$ is a four-dimensional space. Generically, we need to impose four relations among the parameters $a_{i}$ and $b_{j}$ to make $\psi_{0}$ disappear.